D-Night / Eng-Night UT Sep 06 2020

OSA: J. Williams

ISp: C. Bolyard

Mtn Manager: T. Forsyth

SO: J. Power

ISA: B. Rothberg, O. Kuhn

AO: G. Taylor


Summary:

Clouds at sunset. Once they started to clear up we executed an OSCO the completed the White Dwarf D with LUCI-AO ADI. Second half was dedicated to LBC checkout. We had periods of thick cloud cover that forced us to pause, but we completed 3 Superfoc sequences in 0.8-1.1" seeing conditions. We executed a science scripts in the Bessel filters, but did not complete a science script in the remain LBC filters before sunrise. Obtained several LBC skyflats.

Problems:

A few periods of thick clouds. Obs2 experience some slowdowns. RBscience continues to crash with high cadence observations. To start collimation along at the start of the night, we had to apply initial Z4 = -10000nm and Z8 = +2000nm offsets to DX and, to SX, Z7 = +2000. No showstoppers.

Proposed Plan:

  • DONEWD ADI Observations (PI. S.Ertel) - LUCI-AO - (4hrs) (Science)
  • DONELBC Checkout - High Priority
    • DONESuperfocus Sequence (1.5 hrs) - need stable conditions, can tolerate seeing up to 1.2" (if stable)
    • DONE (completed yesterday) Rotator Center Confirmation - (0.5 hrs) - robust agains conditions
    • DONETest Science OB (0.5 hrs) - stable conditions preferred
  • LUCI-AO NS test Eugenia (end of Night) - execute to test system, if </1.2" take full science data (Science)
  • Backups
    • MODS Imaging (possibly spectra) - GRB 200826A (1 hr) - anytime 2nd half night (Science)
    • ESM test - 1more round (M52) - min 0.5hr - depending on discussion with AO team. - backup for poor seeing - exercise AO system, three filters and one AORef mag ~45min with overheads
Nightlog:

00:55 LBC's powered up and taking 25 biases. LUCI's up and Init
  • LBC's 25 Bias
  • LUCI's 10x(1x2.5s) LIR blind luci1.20200906.0001-10. luci2.20200906.0001-10
  • LUCI1 10x(10x2.5s) LIR blind luci1.20200906.0011-20.
  • LUCI1 10x(26x2.5s) LIR blind luci1.20200906.0021-30.
  • LUCI1 10x(6x5s) LIR blind luci1.20200906.0031-40.
1:40 Josh is opening. Verified LUCI1 FS alignment. Greg restarted everything with the WFS and adsec this afternoon.

1:59 Sending preset to OSCO target HD168811. CCD readout fully saturated. Still too bright. There are some clouds tonight.

2:09 Josh is doing a pointing check nerby. Clouds do look like a factor today

2:11 Resending preset. We can't see a 10th magnitude guide star....Clouds too thick, we are in a holding pattern. These clouds just moved in right after we opened.

2:52 Still pretty cloudy but might be clear enough to get some data. Variable extinction...~0-2mag or so seeing 0.9-1.3". OCAM is reporting estimated mag of 9.2-11.2
  • failed Apply optical gain. thick cloud came in why applying optical gain
3:00 resending preset. it does look like we are on the tail end of the really thick stuff.
  • Seeing ~0.75-0.85, estimated mag 8.8 on OCAM
  • luci1.20200906.0043-0062
  • looks like clouds are clearing out.
3:12 Moving to WD1647+591 for collimation, PA 124.

3:16 Loading ADI script. It didn't load the first time. I helps to spell parallactic correctly in my scripts

3:18 Sending ADI script for ADI. We are pointed right at a small cloud.
  • bin1 was probably not going to be robust, so redid acquisition with bin2 to tolerate clouds. Don't need high correction.
  • centered target and absorbed pointing, also tweaked fieldstop. Confirmed rotator is stationary (at 367.7454), Rotator mode parallactic in PCS
  • luci1.20200906.0067
  • 4.2 photons/subapp, ~18 average pupil intensity, bin2 796 Hz, modulation3, estimated 14.1 mag on OCAM. Clouds cleared pretty much right out after intial setup and mostly clear though the majority of the observation, with the ocassional wisp.
  • noted that the images are looking a bit soft luci1.20200906.0175- Greg is reoptimizing. The gains were changing luci1.202009060180-0186 so don't trust those images. The seeing must have deteriorated. the FWHM is larger 6px vs 4px earlier. Image looks much rounder post reoptimization
  • luci1.20200906.0187-0189, safe skipping and opened 0190 looks like we had the gains too high, Reoptimization 0196 with optical gain. Safeskipping right away on 0197. Optical gain is not stable in bin2. 0197-0201 correction not optimal or was being tweaked by greg. Greg manually tweaked and we are back up and running
  • luci1.20200906.202-206, ~ image 206, 2louds in pupil are dropping. looks like we have some clouds coming in, OCAM mag has dropped to 14.8. Lost light by image 209-216. Greg is tweaking the frequency to see if we can get a more stable correction. It drove off of the ccd47
5:40 Resending Preset - image 0217-0218 are setup centering the target, bin2, frequency 757, modulation3, A few cloud wisps moving through and seeing does not looks great. Optimize gain failed so Greg is manually optimizing
  • back on and resuming offset from where we left off: OOPS i did not absorb the pointing offset so target is not in the field for luci1.20200906.0219-0222
  • luci1.20200906.0223- Correction is a bit soft but suspect the seeing has deterioratied and that we are just in the muck.
  • The seeing is definitely variable. The pixel width of our source has varied 4-8.5px and the slope rms on the wfs also suggests seeing is highly variable
  • 0264 safeskipping, reclosed right away.
  • 0268 images are looking quite soft at 9px and elongated now. We are at 32 deg elevation now
  • 0270 safe skipping and opened, reclosed immediately
  • 0277 safeskipping and opened, reclosed immediately
6:45 WD is at 30 deg. Moving up to zenith to reconfigure for LBC's.
Josh is preemtively appling the following to DX Z4 -10000, Z8 +2000

7:13 Slewing 22h36 for collimation and copointing. This field is higher a bit high for Superfoc, we are slewing to a secondary field

7:21 slewing to RA23h25m_DEC48d00m for collimation and pointing. running dohybrid.
  • runaway on blue,
  • clearing optics for both, taking an RB_rVlowextra to see what the pupils look like before we start do hybrid again
  • Adding Z7+2000 based on pupil images. Olga is changing dohybrid coma_thresh to 0.1 to force WRS to only send coma
  • Running dohybrid again. converged with 0.92" seeing
8:02 starting Extra/Intra focal script

8:05 Copointing, I went to a nearby field but not the same one.

8:15 Collimating again. RA23h25m_DEC48d00m
  • it seemed to hang on waiting for image in /newdata. quit and restarted dofpia and converged with 1.05" seeing
8:26 Copointing

8:33 RA23h25m_DEC48d00m El71
  • Running RB_rVexin4
  • Superfoc (mod) RB_rV_sf7 (PA0) RB_rV_sf7_pa180 , RB_rV_sf7 (PA0)
  • HIgh cadence observations RB_science keeps disappearing
  • A few clouds moving in
  • RB_rvexin4
9:04 lots of clouds moving in, moving to RA22h36d
  • dofpia - clouds are too thick. This band of clouds look like they are just passing through with not much behind so we will just wait and see.
  • 9:28 Clouds are getting thicker
10:11 Clouds are quite thick, still in a holding pattern

10:28 The clouds are starting to clear off.
  • dofpia - staying with our field, RA22h36d, now at elev 44 deg
    • converged quickly
  • running the copointing OB, 103527, 103525 ...
  • Ex/In focal pair
    • 103941,103937 and then 104108, 104057
  • superfoc pa=0 (no longer using mod - best focus was seen to be closer to nominal in earlier sequence)
  • 10:50 superfoc pa=180 started
  • 10:58 superfoc pa=0 started
  • 11:06 extra/intra pair
11:08 Slewing to M76 field at 01h51d elev 67 deg
  • dofpia, converged in 2 iterations
  • copointing, 111537 and 111543
    • lbcr copointing image shows the horizontal banding
  • Ex/In focal pair
  • 11:21 superfoc pa0 starting
  • 11:28 superfoc pa180
  • 11:35 superfoc pa0
  • 11:42 extra/intra pair
11:46 Science-like observing on M76
  • 11:46 dofpia on the field, converged in 3 iterations
  • quick check on mirror temps: glass 13.4 ambient 13.8
11:52 M76_UBVRIV
  • guider shows 1.5" on blue and 0.8" on red at U, R but 1" and 0.8" at B. Guide star at U is 1.1 vs 0.8 in the red
12:10 M76 UgrFIR
  • got some images, but saturated
12:18 skyflats:
  • g and I
  • preset error, LBCR rotator failed to move to position commanded.
    2020/09/06 12:24:01.685477 >>> preset/expose/save interrupted at preset time, retcode=-1
    85 2020/09/06 12:24:01.685400 rotator system recovered from a problem -1 [src/channel/channel.c:1166]
    84 2020/09/06 12:24:01.685295 ROTATOR failure [src/rotator/rotator.c:1123]
    83 2020/09/06 12:24:01.685280 ROTATOR actual position is -246.674 deg
    82 2020/09/06 12:24:01.685236 ROTATOR

    actual motion has been 0.000 deg instead of 322.449 deg

  • Stopping the OB and then resending it recovered. Note added to IT 4566 (this is also covered in IT 5889 which was closed in favor of logging new instances and work on the issue in 4566).
  • 12:27 olga is taking flats now for g & I
    • I flats are a bit high biut overall a good set of g & I
  • V & V flats 1 or 2 are ok
  • Uspec & z flats, but the sky was too bright

13:18 Taking a series of 25 LBC biases.

-- BarryRothberg - 05 Sep 2020
Topic revision: r8 - 06 Sep 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback